OpenStudy (marigirl):

Need clarification please Topic: Related rates of change- Differentiation show that: \[\frac{ d }{ dx }(\frac{ v^2 }{ 2 })=\frac{ d^2x }{ dt^2 }\] Where x represents the distance travelled in t seconds by a particle moving with velocity v This is what i have so far:

OpenStudy (marigirl):

\[\frac{ dx }{ dt }=v\] Differentiate this with respect to time and get \[\frac{ d^2x }{ dt^2 }=\frac{ dv }{ dt } \]

OpenStudy (marigirl):

we know: \[\frac{ dv }{ dt }=\frac{ dv }{ dx }\times \frac{ dx }{ dt }\]

OpenStudy (marigirl):

which is \[\frac{ dv }{ dt }=\frac{ dv }{ dx } \times v\]

OpenStudy (marigirl):

im lost now

OpenStudy (anonymous):

By the chain rule, the left hand side is \[\frac{d}{dx}\left[\frac{v^2}{2}\right]=\frac{1}{2}\left(2v\frac{dv}{dx}\right)=v\frac{dv}{dx}\]which you showed is equal to \(\dfrac{dv}{dt}\).

OpenStudy (marigirl):

I am not sure how you for the left side of the equation you wrote above

OpenStudy (anonymous):

Do you know what the chain rule is?

OpenStudy (marigirl):

i think the v and t are confusing me because i keep writing x and y. i might try doing it again